OpenStudy (anonymous):

What are the steps in solving 4/(x+1)-3/(x+2)=1?

OpenStudy (lalaly):

\[\frac{4}{x+1}-\frac{3}{x+2}=1\]\[\frac{4(x+2)-3(x+1)}{(x+1)(x+2)}\]

OpenStudy (lalaly):

\[=1\]

OpenStudy (lalaly):

\[\frac{4x+8-3x-3}{(x+1)(x+2)}=1\]

OpenStudy (lalaly):

\[x+5=(x+1)(x+2)\]

OpenStudy (lalaly):

\[x+5=x^2+3x+2\]\[x^2+2x-3=0\]

OpenStudy (lalaly):

\[(x+3)(x-1)=0\] x=-3 x=1